# Few-shot Class-incremental Learning: A Survey

### v2

@@ 4.1 Data-based Approaches
Data-based approaches refer to addressing FSCIL challenges arising from limited or non-reusable data by focusing on the data perspective. Relevant methodologies include data replay and pseudo-data construction.
@@ 4.1.1 Data Replay-based Methods
Catastrophic forgetting often occurs in FSCIL due to the unavailability of original complete training data from previous sessions. Data replay is a direct strategy to mitigate this issue by replaying valuable data while adapting to new sessions. Existing methods include raw replay and generative replay, involving the replay of samples or feature representations.
Raw Replay-based Methods: The raw replay methods address catastrophic forgetting by storing a portion of raw samples from previous sessions in auxiliary memory and replaying it during the learning process of a new session to review previous knowledge. As shown in Fig. 8(a), Kukleva et al. [28] proposed a multi-stage FSCIL method called LCwoF. It first used the Cross-entropy (CE) loss to train the backbone. In the second stage, it employed the KD loss and base-normalized CE loss to jointly supervise learning new classes and preserve the old knowledge. In the final stage, randomly sampled old and new class data were combined for data replay to further calibrate the performance. Differently, Zhu et al. [44] proposed a feature distribution distillation-based method, which stored the same number of old samples as each new class to form a joint set during its learning process. Both the old and new models generated the feature representations for this set. A joint function based on CE loss and KD loss was employed to constrain the new model to generate similar representations as the old model to preserve the old knowledge. However, the performance of raw replay methods is influenced by factors such as auxiliary storage space, sample selection, and quantity, which have not been fully addressed.
@@
Discussion: Data replay is a direct strategy to address catastrophic forgetting in FSCIL. While raw replay methods provide simplicity and convenience, their effectiveness is influenced by factors including auxiliary storage space, the selection and quantity of samples, and the imbalanced distribution of old and new classes. In comparison, generative replay methods exhibit better flexibility and mitigate potential privacy concerns associated with raw replay methods. However, generative replay methods face challenges in continuously generating old samples in the imbalanced and dynamic data stream, generation quality and efficiency, and additional computational costs. These issues require further exploration and research.
@@ 4.1.2 Pseudo Scenarios-based Methods
Contrasting backward-compatible methods such as data replay that tackle catastrophic forgetting, another prevalent FSCIL strategy is the construction of pseudo-incremental scenarios. These scenarios, acting as rehearsal mechanisms in the dynamic and ever-expanding FSCIL data stream, prepare models for actual incremental sessions, ensuring effective performance. These methods primarily fall into two categories: pseudo-class and pseudo-session construction.
Pseudo Class-based Methods: Pseudo-class construction methods aim to generate synthetic classes and their corresponding samples to facilitate FSCIL models preparing for the real incremental classes. Most current studies employ base sessions to develop these pseudo-classes, training the models using pseudo-data and the original data. This strategy promotes forward compatibility in the FSCIL models. As shown in Fig. 8(c), this approach is the forward-compatible FSCIL framework proposed by Zhou et al. [33]. The crux of this framework lies in constraining the real samples during training, enabling them to render their respective categories more compact in the embedding space and reserve some spaces for the constructed virtual categories. In particular, this method promotes the intra-class compactness of the real data and forces the masked features based on the real data to be closed to a pseudo-class. Simultaneously, the framework employs similar techniques to constrain virtual features constructed from a mixture of multiple class features. It ensures the compactness of real categories while reserving some feature space for incremental classes. Similarly, Peng et al. [47] generated pseudo-classes by merging two distinct classes from the base session and augmenting the data using techniques such as random cropping, horizontal flipping, and color jittering in the ALICE framework. It used angular penalty loss commonly used in face recognition for feature extractor training based on the joint set of pseudo and real data. The core idea also involved promoting intra-class compactness and reserving feature space for incremental classes.
@@ 4.2.1 Dynamic Structure-based Methods
Graph-based Methods: Methods based on graph structures utilize graph topological properties to achieve FSCIL. These methods typically use nodes and edges in the graph to describe the similarity or correlation between different classes from various sessions and adjust the graph structure based on the mutual influences among classes. Some studies employ graph structures to implement FSCIL [13, 14]. For example, Tao et al. [13] proposed the TOPIC framework, which utilizes the neural gas network for knowledge extraction and representation. TOPIC aimed to address FSCIL by dynamically adjusting the interrelationships between feature representations. Specifically, the neural gas network defined an undirected graph that mapped the feature space to a finite set of feature vectors and maintained the topological properties of the feature space through competitive Hebbian learning [70]. To achieve FSCIL, they gradually improved the neural gas network by enabling the supervised neural gas model to grow nodes and edges through competitive learning. Additionally, they designed a stability loss to suppress catastrophic forgetting and an adaptability loss to reduce overfitting. In addition, the CEC framework [14] mentioned in Sec. 4.1.2 also utilized graph structures for FSCIL. It first trained a graph attention network using pseudo-incremental sessions to adjust the model. During incremental sessions, the model incorporated an attention mechanism to regulate the interrelationships between nodes, represented by prototypes of old and new classes. This allowed for better context information transfer between sessions, making the class prototypes more robust.
@@
Discussion: Dynamic structure methods are important approaches to address the challenges of FSCIL. These methods achieve the learning of new knowledge while preserving old knowledge by dynamically adjusting the model structure or the relationships between prototypes. For example, graph-based methods utilized the topological characteristics of graphs to achieve non-forgettable IL by adjusting nodes and edges to describe the similarity and correlation between different classes. Dynamic structure networks enhances feature representation and alleviates overfitting by temporarily expanding and dynamically compressing network nodes. Dynamic structural methods play a significant role in FSCIL, but further research and exploration are still needed to develop more design methods for dynamic structures.
@@ 4.2.2 Attention-based Methods
Attention-based methods in FSCIL adjust the attention allocation of features by introducing attention modules into the model structure. This allows the model to focus on information relevant to the current task, improving its performance and generalization ability. The role of attention modules used in many FSCIL approaches [73, 14, 16, 74, 27] is diverse. For example, in the dual-branch KD framework proposed by Zhao et al. [73], which consisted of a base branch and a novel branch, they noted that fine-tuning by novel classes inevitably led to forgetting old knowledge. To further improve the performance of base classes, they proposed an attention-based aggregation module that selectively merges predictions from the base branch and the novel branch. Furthermore, Cheraghian et al. [27] employed meta-learning to train a backbone that can incrementally learn new classes with limited samples without forgetting the old classes. However, many existing FSCIL paradigms update the classifier by concatenating the base classifier with the new class prototypes obtained by averaging the features of each training sample. This approach often leads to bias. Therefore, this paper proposes a correction model based on Transformer [75]. With its attention mechanism, the correction model can effectively transmit context information among different classes, making the classifier more efficient and robust. Similar approaches include the graph attention network used in the CEC framework [14].
@@ 4.3.1 Representation Learning-based Methods
Metric Learning-based Methods: Metric learning aims to determine the similarity between objects using an optimal distance metric for learning tasks [77]. It has found extensive application in FSL [78]. Recently, metric learning has also been adopted in FSCIL to learn effective representations. Among the commonly used approaches, triplet loss stands out. As shown in Fig. 9(a), Mazumder et al. [31] proposed a novel approach for FSCIL. It incorporated self-supervised learning to enhance the generalization capability of the backbone. Then, this approach analyzed the importance of model parameters and updated only the unimportant parameters for new classes. The update was achieved by combining three loss functions: triplet loss, regularization loss, and cosine similarity loss. The triplet loss aimed to generate discriminative features, while the regularization loss prevented catastrophic forgetting. The cosine similarity loss focuses on controlling the similarity between old and new prototypes. Thus, effective performance is achieved for FSCIL. Moreover, other metric learning methods are applied to FSCIL too. Concretely, Peng et al. [47] proposed the ALICE framework, incorporating the angular penalty loss originally used in face recognition to obtain well-clustered features. This loss was employed to train the backbone using both base class data and synthetic data, thereby creating additional space for accommodating incremental classes, and cosine similarity was applied to achieve the classification.
Figure 9: The representative classification frameworks in FSCIL: (a) FSLL+SS, using self-supervised learning and metric learning to learn the backbone and evaluate unimportant parameters for new class learning; (b) NC-FSCIL, applying neural collapse theory to pre-define a classifier for backbone learning; (c) FeSSSS, merging features from self-supervised and supervised learning; and (d) Us-KD, leveraging unlabeled data with uncertainty quantification to perform KD.
@@
Despite the good performance achieved by these margin-based metric losses, Zou et al. [36] pointed out the issue in FSCIL: large margin values can result in good discriminability among base classes but hinder the generalization capability of new classes. Conversely, small or even negative margin values can lead to poor performance on the base classes but exhibit better generalization on new classes. This phenomenon is known as the class-level overfitting problem. To address this issue, Zou et al. [36] proposed the CLOM framework, which combined margin theory with the characteristics of neural network structures. Specifically, since the shallow layers of neural networks are more suitable for learning common features among classes, while the deep layers are better suited to acquiring advanced features, they designed a loss function that constrains shallow feature learning and deep feature learning separately. Furthermore, this framework alleviates the class-level overfitting problem by integrating class relationships.
@@
Furthermore, some methods address FSCIL by designing feature spaces of specialized structures. For example, Hersche et al. [34] proposed a C-FSCIL framework, which consisted of a feature extractor trained by meta-learning, a trainable fully connected layer, and a rewritable explicit memory. The core idea was introducing hyperdimensional embedding, which has three advantages: the high probability of quasi-orthogonality between random vectors, (2) rich expressive space, and (3) good semantic representation capability. C-FSCIL had three training strategies. The first was based on simple meta-learning, as described in Sec. 4.3.3. The second strategy stored initial prototypes in the globally average activation memory and applied an element-wise sign operation to transform similar feature prototypes into bipolar vectors. These transformed vectors were then supervised to train the fully connected layer, which learned the weights for the final prototypes. The third strategy was similar to the second one but incorporated two losses to constrain the inter-class differences and maintain the relationship with the original prototypes. Besides, as shown in Fig. 9(b), Yang et al. [82] proposed an FSCIL framework based on neural collapse [83], which refers to the phenomenon that at the end of training (after 0 training error rate), the last layer features of the same class collapse into a single vertex in the feature space, aligning all class vertices with their classifier prototypes and forming as a simplex quiangular tight frame. Based on this characteristic, the proposed framework predefined a structure similar to neural collapse and directed the model to optimize it. Specifically, a group of prototypes for both the base and incremental sessions was pre-assigned as a simplified form of quiangular tight frame. During training, the prototypes were fixed. They introduced a novel loss function and an additional projection layer to assign each class to its respective prototype separately. Without cumbersome operations, this method achieved superior performance. In addition, in Sec. 4.1.2, the method proposed by Zhou et al. [33] also addressed FSCIL by learning the feature space in a way that preserves some space for incremental classes during the learning of base classes.
@@
In addition to feature fusion based on self-supervised features, there are also works that integrate other features to achieve good performance in FSCIL. For example, Yao et al. [90] proposed a simple strategy for enhancing the new prototype. Specifically, it first trained a CNN network on the base classes and kept it fixed. It was used to generate class prototypes for both base and new classes. Then, the initial class prototypes for new classes were measured for their similarity to base class prototypes. Based on the similarity, the prototypes for new classes were updated by mixing the initial weights with other base prototypes. This fusion enhancement strategy imitated human cognition by guiding new class learning based on existing knowledge.
@@ 4.3.2 Knowledge Distillation-based Methods
Another solution to mitigate data imbalance and limited samples in FSCIL is introducing additional data to prevent overfitting. In the context of FSCIL, Cui et al. proposed a series of semi-supervised methods that leverage KD and unlabeled data [93, 45, 46]. As shown in Fig. 9(d), Cui et al. [45] introduced the Us-KD framework, which used an uncertainty-guided module to select unlabeled data to mitigate overfitting during knowledge transfer. The framework initially trained the model on base classes and stored some samples. In the incremental session, the model was initialized with the previous model’s weights and updated using stored old samples and labeled samples from the current session through classification and distillation losses. Subsequently, the uncertainty-guided module selected and labeled unlabeled samples, combined with labeled samples to update the model iteratively. Finally, the stored data was updated with these samples. In their further research [46], they pointed out that well-learned or easily classifiable classes often have higher prediction probabilities. Thus, they designed a data selection method called ”Class Equilibrium,” where well-learned categories are assigned fewer samples, and poorly learned categories are assigned more samples. It is worth noting that they highlighted the potential unreliability of KD with unlabeled data. Thus, they introduced an uncertainty-aware distillation approach suitable for semi-supervised FSCIL, consisting of uncertainty-guided refinement and adaptive distillation loss. Refinement involved leveraging uncertainty assessment to filter reliable samples from the augmented dataset, while adaptive distillation adjusted the distillation loss weights based on the sample quantity.

### v3

@@ 4.1 Data-based Approaches
Data-based approaches refer to addressing FSCIL challenges arising from limited or non-reusable data by focusing on the data perspective. Relevant methodologies include DR and pseudo-data construction.
@@ 4.1.1 Data Replay-based Methods
Catastrophic forgetting often occurs in FSCIL due to the unavailability of original complete training data from previous sessions. DR is a direct strategy to mitigate this issue by replaying valuable data while adapting to new sessions. Existing methods include raw replay and generative replay, involving the replay of samples or feature representations.
Raw Replay-based Methods: The raw replay methods address catastrophic forgetting by storing a portion of raw samples from previous sessions in auxiliary memory and replaying it during the learning process of a new session to review previous knowledge. As shown in Fig. 8(a), Kukleva et al. [18] proposed a multi-stage FSCIL method called LCwoF. It first used the Cross-entropy (CE) loss to train the backbone. In the second stage, it employed the KD loss and base-normalized CE loss to jointly supervise learning new classes and preserve the old knowledge. In the final stage, randomly sampled old and new class data were combined for DR to further calibrate the performance. Differently, Zhu et al. [44] proposed a feature distribution distillation-based method, which stored the same number of old samples as each new class to form a joint set during its learning process. Both the old and new models generated the feature representations for this set. A joint function based on CE loss and KD loss was employed to constrain the new model to generate similar representations as the old model to preserve the old knowledge. However, the performance of raw replay methods is influenced by factors such as auxiliary storage space, sample selection, and quantity, which have not been fully addressed.
@@
Discussion: DR is a direct strategy to address catastrophic forgetting in FSCIL. While raw replay methods provide simplicity and convenience, their effectiveness is influenced by factors including auxiliary storage space, the selection and quantity of samples, and the imbalanced distribution of old and new classes. In comparison, generative replay methods exhibit better flexibility and mitigate potential privacy concerns associated with raw replay methods. However, generative replay methods face challenges in continuously generating old samples in the imbalanced and dynamic data stream, generation quality and efficiency, and additional computational costs. These issues require further exploration and research.
@@ 4.1.2 Pseudo Scenarios-based Methods
Contrasting backward-compatible methods such as DR that tackle catastrophic forgetting, another prevalent FSCIL strategy is the construction of pseudo-incremental scenarios. These scenarios, acting as preview mechanisms in the dynamic and ever-expanding FSCIL data stream, prepare models for actual incremental sessions, ensuring effective performance. These methods primarily fall into two categories: pseudo-class and pseudo-session construction.
Pseudo Class-based Methods: Pseudo-class construction methods aim to generate synthetic classes and their corresponding samples to facilitate FSCIL models preparing for the real incremental classes. Most current studies employ base sessions to develop these pseudo-classes, training the models using pseudo-data and the original data. This strategy promotes forward compatibility in the FSCIL models. As shown in Fig. 8(c), this approach is the forward-compatible FSCIL framework proposed by Zhou et al. [23]. The crux of this framework lay in constraining the real samples during training, enabling them to render their respective categories more compact in the embedding space and reserve some spaces for the constructed virtual categories. In particular, this method promoted the intra-class compactness of the real data and forced the masked features based on the real data to be closed to a pseudo-class. Simultaneously, the framework employed similar techniques to constrain virtual features constructed from a mixture of multiple class features. It ensured the compactness of real categories while reserving some feature space for incremental classes. Similarly, Peng et al. [47] generated pseudo-classes by merging two distinct classes from the base session and augmenting the data using techniques such as random cropping, horizontal flipping, and color jittering in the ALICE framework. It used angular penalty loss commonly used in face recognition for feature extractor training based on the joint set of pseudo and real data. The core idea also involved promoting intra-class compactness and reserving feature space for incremental classes.
@@ 4.2.1 Dynamic Structure-based Methods
Graph-based Methods: Methods based on graph structures utilize graph topological properties to achieve FSCIL. These methods typically use nodes and edges in the graph to describe the similarity or correlation between different classes from various sessions and adjust the graph structure based on the mutual influences among classes. Some studies employ graph structures to implement FSCIL [13, 14]. For example, Tao et al. [13] proposed the TOPIC framework, which utilized the neural gas network for knowledge extraction and representation. TOPIC aimed to address FSCIL by dynamically adjusting the interrelationships between feature representations. Specifically, the neural gas network defined an undirected graph that mapped the feature space to a finite set of feature vectors and maintained the topological properties of the feature space through competitive Hebbian learning [70]. To achieve FSCIL, they gradually improved the neural gas network by enabling the supervised neural gas model to grow nodes and edges through competitive learning. Additionally, they designed a stability loss to suppress catastrophic forgetting and an adaptability loss to reduce overfitting. In addition, the CEC framework [14] mentioned in Sec. 4.1.2 also utilized graph structures for FSCIL. It first trained a graph attention network using pseudo-incremental sessions to adjust the model. During incremental sessions, the model incorporated an attention mechanism to regulate the interrelationships between nodes, represented by prototypes of old and new classes. This allowed for better context information transfer between sessions, making the class prototypes more robust.
@@
Discussion: Dynamic structure methods are important approaches to address the challenges of FSCIL. These methods achieve the learning of new knowledge while preserving old knowledge by dynamically adjusting the model structure or the relationships between prototypes. For example, graph-based methods utilize the topological characteristics of graphs to achieve non-forgettable IL by adjusting nodes and edges to describe the similarity and correlation between different classes. Dynamic structure networks enhances feature representation and alleviates overfitting by temporarily expanding and dynamically compressing network nodes. Dynamic structural methods play a significant role in FSCIL, but further research and exploration are still needed to develop more design methods for dynamic structures.
@@ 4.2.2 Attention-based Methods
Attention-based methods in FSCIL adjust the attention allocation of features by introducing attention modules into the model structure. This allows the model to focus on information relevant to the current task, improving its performance and generalization ability. The role of attention modules used in many FSCIL approaches [73, 14, 16, 74, 17] is diverse. For example, in the dual-branch KD framework proposed by Zhao et al. [73], which consisted of a base branch and a novel branch, they noted that fine-tuning by novel classes inevitably led to forgetting old knowledge. To further improve the performance of base classes, they proposed an attention-based aggregation module that selectively merges predictions from the base branch and the novel branch. Furthermore, Cheraghian et al. [17] employed meta-learning to train a backbone that can incrementally learn new classes with limited samples without forgetting the old classes. However, many existing FSCIL paradigms updated the classifier by concatenating the base classifier with the new class prototypes obtained by averaging the features of each training sample. This approach often led to bias. Therefore, this paper proposed a correction model based on Transformer [75]. With its attention mechanism, the correction model can effectively transmit context information among different classes, making the classifier more efficient and robust. Similar approaches included the graph attention network used in the CEC framework [14].
@@ 4.3.1 Representation Learning-based Methods
Metric Learning-based Methods: Metric learning aims to determine the similarity between objects using an optimal distance metric for learning tasks [77]. It has found extensive application in FSL [78]. Recently, metric learning has also been adopted in FSCIL to learn effective representations. Among the commonly used approaches, triplet loss stands out. As shown in Fig. 9(a), Mazumder et al. [21] proposed a novel approach for FSCIL. It incorporated self-supervised learning to enhance the generalization capability of the backbone. Then, this approach analyzed the importance of model parameters and updated only the unimportant parameters for new classes. The update was achieved by combining three loss functions: triplet loss, regularization loss, and cosine similarity loss. The triplet loss aimed to generate discriminative features, while the regularization loss prevented catastrophic forgetting. The cosine similarity loss focused on controlling the similarity between old and new prototypes. Thus, FSCIL achieved effective performance. Moreover, other metric learning methods are applied to FSCIL too. Concretely, Peng et al. [47] proposed the ALICE framework, incorporating the angular penalty loss originally used in face recognition to obtain well-clustered features. This loss was employed to train the backbone using both base class data and synthetic data, thereby creating additional space for accommodating incremental classes, and cosine similarity was applied to achieve the classification.
Figure 9: The representative classification methods in FSCIL, with the core designs highlighted in red. (a) FSLL+SS initially uses self-supervised learning to provide good generalization. Then, it evaluates the importance of model parameters, fixing important ones to maintain old knowledge while using metric learning to learn new classes with unimportant parameters; (b) NC-FSCIL uses neural collapse theory to pre-define a classifier before training, and the model is constrained to learn towards the pre-defined classifier via the projection layer, preventing conflicts between learning new and old classes; (c) FeSSSS introduces a ResNet50 trained through self-supervised learning in addition to the regular ResNet backbone. By merging the supervised and self-supervised features, the framework enhances the generalization capability in FSCIL; (d) Us-KD focuses on semi-supervised FSCIL, first utilizing KD with stored old samples and new data to learn new knowledge while retaining old knowledge. Then, uncertainty quantification is applied to select suitable unlabeled data for labeling, combined with labeled samples for iterative model updates.
@@
Despite the good performance achieved by these margin-based metric losses, Zou et al. [26] pointed out the issue in FSCIL: large margin values can result in good discriminability among base classes but hinder the generalization capability of new classes. Conversely, small or even negative margin values can lead to poor performance on the base classes but exhibit better generalization on new classes. This phenomenon is known as the class-level overfitting problem. To address this issue, Zou et al. [26] proposed the CLOM framework, which combined margin theory with the characteristics of neural network structures. Specifically, since the shallow layers of neural networks are more suitable for learning common features among classes, while the deep layers are better suited to acquiring advanced features, they designed a loss function that constrains shallow feature learning and deep feature learning separately. Furthermore, this framework alleviated the class-level overfitting problem by integrating class relationships.
@@
Furthermore, some methods address FSCIL by designing feature spaces of specialized structures. For example, Hersche et al. [24] proposed a C-FSCIL framework, which consisted of a feature extractor trained by meta-learning, a trainable fully connected layer, and a rewritable explicit memory. The core idea was introducing hyperdimensional embedding, which has three advantages: (1) the high probability of quasi-orthogonality between random vectors, (2) rich expressive space, and (3) good semantic representation capability. C-FSCIL had three training strategies. The first was based on simple meta-learning, as described in Sec. 4.3.3. The second strategy stored initial prototypes in the globally average activation memory and applied an element-wise sign operation to transform similar feature prototypes into bipolar vectors. These transformed vectors were then supervised to train the fully connected layer, which learned the weights for the final prototypes. The third strategy was similar to the second one but incorporated two losses to constrain the inter-class differences and maintain the relationship with the original prototypes. Besides, as shown in Fig. 9(b), Yang et al. [82] proposed an FSCIL framework based on neural collapse [83], which refers to the phenomenon that at the end of training (after 0 training error rate), the last layer features of the same class collapse into a single vertex in the feature space, aligning all class vertices with their classifier prototypes and forming as a simplex equiangular tight frame. Based on this characteristic, the proposed framework predefined a structure similar to neural collapse and directed the model to optimize it. Specifically, a group of prototypes for both the base and incremental sessions was pre-assigned as a simplified form of equiangular tight frame. During training, the prototypes were fixed. They introduced a novel loss function and an additional projection layer to assign each class to its respective prototype separately. Without cumbersome operations, this method achieved superior performance. In addition, in Sec. 4.1.2, the method proposed by Zhou et al. [23] also addressed FSCIL by learning the feature space in a way that preserves some space for incremental classes during the learning of base classes.
@@
In addition to feature fusion based on self-supervised features, there are also works that integrate other features to achieve good performance in FSCIL. For example, Yao et al. [90] proposed a simple strategy for enhancing the new prototype. Specifically, it first trained a Convolutional Neural Network (CNN) on the base classes and kept it fixed. It was used to generate class prototypes for both base and new classes. Then, the initial class prototypes for new classes were measured for their similarity to base class prototypes. Based on the similarity, the prototypes for new classes were updated by mixing the initial weights with other base prototypes. This fusion enhancement strategy imitated human cognition by guiding new class learning based on existing knowledge.
@@ 4.3.2 Knowledge Distillation-based Methods
Another solution to mitigate data imbalance and limited samples in FSCIL is introducing additional data to prevent overfitting. In the context of FSCIL, Cui et al. proposed a series of semi-supervised methods that leverage KD and unlabeled data [93, 45, 46]. As shown in Fig. 9(d), Cui et al. [45] introduced the Us-KD framework, which used an uncertainty-guided module to select unlabeled data to mitigate overfitting during knowledge transfer. The framework initially trained the model on base classes and stored some samples. In the incremental session, the model was initialized with the previous model’s weights and updated using stored old samples and labeled samples from the current session through classification and distillation losses. Subsequently, the uncertainty-guided module selected and labeled unlabeled samples, combined with labeled samples to update the model iteratively. Finally, the stored data was updated with these samples. In their further research [46], they pointed out that well-learned or easily classifiable classes often have higher prediction probabilities. Thus, they designed a data selection method called ”Class Equilibrium,” where well-learned categories were assigned fewer samples, and poorly learned categories were assigned more samples. It is worth noting that they highlighted the potential unreliability of KD with unlabeled data. Thus, they introduced an uncertainty-aware distillation approach suitable for semi-supervised FSCIL, consisting of uncertainty-guided refinement and adaptive distillation loss. Refinement involved leveraging uncertainty assessment to filter reliable samples from the augmented dataset, while adaptive distillation adjusted the distillation loss weights based on the sample quantity.